# Masked Mixers for Language Generation and Retrieval

### v1

@@ 5 Limitations of this work
This study has been limited in scale: experiments were performed using a relatively small dataset and compute quantity. Although there are reasons to suspect that the results here would translate to much larger datasets and compute such as better scaling properties to larger context windows during training for masked mixers compared to transformers, this has yet to be explored. Another limitation inherent in this work is that compute efficiency is measured in a concrete hardware-specific fashion. We chose this route as it is immediately applicable to the primary goal of language model training (ie getting the lowest possible loss in the least amount of time), and one can recover the FLOPS per experiment using available data on each compute node.
A downside to our hardware-specific approach is that our GPUs (V100s and RTX 3060) are not typically used for language model training in industrial settings today, where instead H100 and A100s are favored. Newer GPUs are specially tailored for transformer inference and training and thus our comparisons between transformer and mixer architectures would likely be somewhat different if performed on these newer GPUs. H100s in particular have dedicated transformer engines [26], which would be expected to provide speedups for that model relative to others.
It is also unclear how masked mixers would perform if given datasets with much more varied samples. TinyStories has a very large number of very similar data samples, which makes learning language generation relatively easy compared to more natural datasets used to train much larger models. On the other hand, the similarities in the inputs results in much more difficult for models to learn to match story summaries to stories: for a batch of 128 stories, there are typically dozens about a main character named ‘Tim’ who ends up having very similar exploits.
@@ 6 Discussion
6.1 Input representation Accuracy and model efficiency
This work has explored the use of input representation accuracy as a guide to improving deep learning model architectures for language generation and retrieval. Despite this perspective allowing for the improvement of model efficiency for both tasks, it has also become is evident that there is no simple relationship between representation accuracy and either generation or retrieval or else very small masked mixers would outperform transformers, and untrained mixers would for effective embeddings for retrieval. Instead it is evident that other considerations (for example, the set of possible transformations a model is capable) are important in varying degrees.
Accurate input representation is evidently useful for creating language embeddings, but one might wonder if this property would also contribute to a model’s ability to memorize the inputs it was trained on. In other words, if a model is capable of retaining all the information in its input then is it prone to memorization? In this work we found that indeed the gap between training and test loss is slightly higher for masked mixers compared to transformers and in that sense the answer is yes. But we have yet to see any examples of severe overfitting of a language dataset of significant size for any masked mixer or transformer, which provides support for the ideas that memorization as it is currently observed is often a property of the dataset itself rather than a model’s architecture and that even very large models capable of memorizing large datasets are intrinsically biased against doing so when training using gradient descent due to a few characteristics of high-dimensional space [27].
6.2 Recurrent, semi-recurrent, and feedforward architectures
@@
This lack of recurrent-like character for the masked mixer results in slightly longer evaluation or text generation times than for transformers of equivalent size. Language retrieval text corpora are typically chunked such that every input is approximately the same size, which means that transformers and mixers with the same number of operations would have very similar runtimes for the task of language embedding. That said, it is not difficult to imagine ways to reduce the memory and space requirements for masked mixers during inference: this may be done via lazy loading of convolutional weights only when they are required, or else by caching previously computed values.
@@ 6.3 Is attention useful for efficient language modeling?
The MLP mixer was originally introduced to test whether or not attention is all you ‘need’, and the answer was found to be no: in contrast, swapping feedforward MLP layers with attention in a vision transformer led to near-complete failures during training indicating that self-attention is not as important at least as the feedforward layers [11]. In this work we have found that attention is not necessary for training an effective generative language model as well, but primarily focus on the efficiency of the training process rather than its asymptotic characteristics.
@@
In this work we have seen evidence both for and against the idea that attention mechanisms are important for learning efficiency: the best attention-only models are slightly more efficient learners than convolution-only models, but on the other hand these attention models have been extensively optimized relative to the masked mixers introduced here. Given a similar effort to optimize the convolutional operations in the mixer, it is possible that this architecture would be more efficient than the most recent transformers.
From the results presented in this paper it would appear that properly optimized attention mechanisms do confer substantial benefits to the process of causal language model training, at least until more efficient training can be obtained using attentionless models such as masked mixers. On the other hand the evidence suggests that attention is generally not desirable for efficient language retrieval tasks, either in the generative model creating an embedding of each input or in a retrieval model trained to match embeddings.
@@ 6.4 How many parameters are necessary for language modeling?
For more concrete estimates based on datasets used today, consider that language models today are nearly universally trained to predict every ‘next’ token in a concatenated sequence of tokens. This means that for a training dataset of $n$ tokens, the task is approximate $n$ points in some presumably very high-dimensional space. The largest dataset any open-source model has been trained on is 15 trillion tokens (for Llama-3 [5]), and in this case we can represent this dataset arbitrarily well in a space of approximately $2408\ln(15\times 10^{12})\approx 8*30=$ 240 dimensional. The goal of language models is to generalize to larger datasets and thus we would hope the model would accurately predict the next tokens of a much larger dataset. But even supposing that this 15 trillion token dataset is only one millionth of the size of this generalized dataset, one would still only require a space of $3528\ln(15\times 10^{18})\approx 8*44=$ 352 dimensions.
How does this relate to the number of parameters necessary in a model? Assuming that the training algorithm is sufficiently powerful (which is not at all a safe assumption, more on this later) the number of parameters in a model could correspond to one of two things: either it is equivalent to the dimensionality of the model with respect to the points it can approximate, or else the model’s ’width’ or hidden layer dimension is equivalent to the model’s dimensionality. Because the vector space of successive layers of practically every modern deep learning model are dependent (activations in layer $n+1$ depend on layer $n$ and perhaps layer $n-1$ etc.) and dimensionality is defined on linear independence, it seems more likely that a model’s dimensionality best corresponds with its hidden layer dimension. If this is true then a model of no more than around 1300 hidden layer elements should be capable of completing any English language sentence, or a model with a width of 350 can accurately predict a next token for a massive dataset. If these hidden widths were used for default Llama architectures, the resulting models would are around 100 million (given a hidden width of $d_{m}=350$) and 995 million (if $d_{m}=1300$) parameters.
It cannot necessarily be assumed that a model with that number of parameters is actually trainable, however: it could be that training requires a large model that must then be converted into a small model. This is the approach used when performing pruning, where parameters are dropped depending on their importance for some output. Alternatively, instead of removing parameters one could reduce the memory required to store each parameters: this is the approach of quantization methods, which are perhaps the most effective methods currently available for shrinking the effective size of a model. The observation that weight quantization rather than pruning is the most effective method for reducing a transformer model’s effective size suggests that this particular architecture may indeed require nearly all the trained parameters in order to function effectively, although whether this is the case or not remains an open questions.

### v2

@@ 7 Limitations of this work
This study is limited in scale: experiments were performed using a relatively small datasets (none larger than 10 billion tokens) and compute (four V100 GPUs) and model sizes (tens or hundreds of millions of parameters). Although there are reasons to suspect that the causal language modeling results here would translate to much larger datasets and compute, this has yet to be investigated as training runs were performed in 30h or less on 4x V100s (for a maximum theoretical limit of $5.4*10^{19}$ Tensor FLOPs), which is a relatively small amount by modern standards. Another limitation inherent in this work is that compute efficiency is measured in a concrete hardware-specific fashion. We chose this route as it is immediately applicable to the primary goal of language model training (ie getting the lowest possible loss in the least amount of time), and one can recover the FLOPS per experiment using available data on each compute node.
A downside to our wall clock and hardware-specific performance measurements is that our GPUs are not typically used for language model training in industrial settings today, where H100 or A100s are favored. Newer GPUs are specially tailored for transformer inference and training and thus our comparisons between transformer and mixer architectures would likely be somewhat different if performed on these newer GPUs. H100s in particular have dedicated transformer engines [34], which would be expected to provide speedups for that model relative to other architectures.
Our input representation methods yield results that are consistent with other invertibility investigations (autoencoding performance for example) but are by no means comprehensive. It is entirely possible that a different input representation method would give more accuracy to trained transformers or less to masked mixers.
@@ 8 Discussion
8.1 Input representation Accuracy and Model Efficiency: What Good is Invertiblity for Deep Learning?
This work has explored the use of input representation accuracy as a guide to improving deep learning model architectures for language generation and retrieval, and has found that global invertibility is indeed very useful in estimating performance capabilities. This is not at all a surprising result when viewed from the theory of numerical analysis, as invertibility is usually the second-most important facet of a system (after linearity). It has also become is evident that there is no simple relationship between representation accuracy and either generation or retrieval or else very small masked mixers would outperform transformers, and untrained mixers would for effective embeddings for retrieval. It is evident that other considerations (for example, the set of possible transformations a model is capable) are important in varying degrees as well.
One might wonder if accurate input representation would contribute to a model’s ability to overfit the inputs it was trained on. In this work we found that indeed the gap between training and test loss is slightly higher for masked mixers compared to transformers and in that sense the answer is yes. But we have yet to see any examples of severe overfitting of a language dataset of significant size for any masked mixer or transformer, which provides support for the ideas that memorization as it is currently observed is often a property of the dataset itself rather than a model’s architecture and that even very large models capable of memorizing large datasets are intrinsically biased against doing so when training using gradient descent due to a few characteristics of high-dimensional space [35].
8.2 Semi-recurrent and feedforward architectures
@@
We can also comment on similarities and differences between the masked mixer and other language models replacing attention with specialized convolutions such as the Monarch mixer [14] or Hyena [15] architectures, which are motivated primarily by an effort to reduce the $O(n^{2}d)$ computational complexity inherent in the transformer architecture. Firstly, we note that although all three works are focused on increasing the efficiency of language modeling, our work has an entirely different motivation (increased representational efficiency rather than lower bounds for computational scaling) and thus focuses on training efficiency at the expense of inference efficiency. In particular, masked mixers have higher inference complexity ($O(n^{3}d)$) than transformers, which can cache hidden layer values. Caching masked mixer hidden state values during inference would be useful for convolutions with restricted substructure (for example, parameterized as Toeplitz or Monarch matrices) but we leave the investigation of such models to future work.
A major difference between this work and many comparisons of sequence-length efficient architectures to transformers in the literature is that the latter do not typically use transformers with the latest architectural features combined with extensive hyperparameter and model size optimizations when establishing a performance baseline for a GPT-style causal language model. There is often good reason for this: transformers have benefited from years of optimizations that often drastically improve training efficiency and new architectures have not had this kind of effort applied.
8.3 How much more efficient are masked mixers compared to transformers for retrieval?
The results in Section 6.3 indicate that masked mixers are substantially more effective retrieval models compared to transformers receiving the same training protocol, and we include results from the near-SOTA e5 Mistral instruct as a benchmark of what a very capable off-the-shelf model can do for this dataset. Finding exactly how much more efficient the masked mixer architecture is for retrieval would require knowledge of the scaling properties of each model type, work that we leave to future investigations.
That said, it is helpful to note the scale of the computational resources with which these masked mixers were trained relative to e5 Mistral. It is unknown exactly how many GPUs and for how long the Mistral 7b model was pretrained, or if this model was indeed trained from scratch without weight transfer, but interpolating this model’s performance with others for which data is available such as Llama 2 (8b) which required 184k A100 GPU-hours and Llama 3.1 (8b) which required 1.46M H100 hours suggests that the 7b Mistral model would require perhaps 500k A100 GPU-hours for pretraining, as a rough ballpark figure. Assuming an equivalence of 1.5 * V100 = A100 (which is empirically a fairly accurate value for for 16/32 bit mixed precision training) we get 750k V100 GPU-hours, which is approximately 10,000x the amount of compute that the transformer and masked mixer was pretrained with (80 GPU-hours). The e5 Mistral Instruct model was then retrieval-trained with 32 V100s over days, which is between 100x and 1,000x the compute we used here for our retrieval training process.
With this in mind we do not mean to imply that masked mixers are a factor of 10,000 more efficient than transformers for retrieval, only that a result achieved using a large transformer model may be surpassed with a small masked mixer using 1/10,000th the compute. We present evidence for the hypothesis that this results primarily from the use of masked mixers for retrieval and also from the training of this model on I.I.D. data to the evaluation set.
@@ 8.4 Is attention useful for efficient language modeling?
The MLP mixer was originally introduced to test the importance of attention in the transformer architecture’s effective modeling, and it was found that swapping attention for MLPs does not lead to a large drop in vision modeling performance but that swapping MLPs for attention does [11]. In this work we have found that attention is not necessary for training an effective generative language model as well, but primarily focus on the efficiency of the training process rather than its asymptotic characteristics.
@@
From the results presented in this paper it would appear that properly optimized attention mechanisms do confer benefits to the process of efficient causal language model training for larger context windows, at least until more efficient training can be obtained using attentionless models such as masked mixers. On the other hand the evidence suggests that attention is generally not desirable for language retrieval tasks, as is predicted by the model invertibility framework presented in this paper.
@@ 8.5 How many parameters are necessary for language modeling?
For more concrete estimates based on datasets used today, consider that language models today are nearly universally trained to predict every ‘next’ token in a concatenated sequence of tokens. This means that for a training dataset of $n$ tokens, the task is approximate $n$ points in some presumably very high-dimensional space. The FineWeb-edu 10BT dataset would thus require a model of dimension $1848\ln(10\times 10^{9})\approx 8*23=$ 184 for perfect memorization, and this number grows extremely slowly as the dataset size increases. The largest dataset open-source models are currently trained on is 15 trillion tokens (for Llama-3 [5]), and in this case we can represent this dataset arbitrarily well in a space of approximately $2408\ln(15\times 10^{12})\approx 8*30=$ 240 dimensional. The goal of language models is to generalize to larger datasets and thus we would hope the model would accurately predict the next tokens of a much larger dataset. But even supposing that this 15 trillion token dataset is only one millionth of the size of this generalized dataset, one would still only require a space of dimension $3528\ln(15\times 10^{18})\approx 8*44=$ 352.
How does this relate to the number of parameters necessary in a model? Two options are either it is equivalent to the dimensionality of the model with respect to the points it can approximate, or else the model’s ‘width’ or hidden layer dimension is equivalent to the model’s dimensionality. Because the vector space of successive layers of practically every modern deep learning model are dependent (activations in layer $n+1$ depend on layer $n$ and perhaps layer $n-1$ etc.) and dimensionality is defined on linear independence, it seems more likely that a model’s dimensionality best corresponds with its hidden layer dimension. If this is true then a model of no more than around 1300 hidden layer elements should be capable of completing any English language sentence, or a model with a width of 350 can accurately predict a next token for a massive dataset. If these hidden widths were used for default Llama architectures, the resulting models would are around 100 million (given a hidden width of $d_{m}=350$) and 995 million (for $d_{m}=1300$) parameters.